# Optimism Without Regularization: Constant Regret in Zero-Sum Games

### v1

@@ 4.1 Subspace Dynamics of Optimistic Fictitious Play
$P_{2}$ =$\;=\;\{z\in\mathbb{R}^{2}:z_{1}<0\;\text{and}\;z_{2}>0\}\qquad P_{3}\;=\;\{z\in\mathbb{R}^{2}:z_{1}>0\;\text{and}\;z_{2}>0\}$ (7)
$P_{1}$ =$\;=\;\{z\in\mathbb{R}^{2}:z_{1}<0\;\text{and}\;z_{2}<0\}\qquad P_{4}\;=\;\{z\in\mathbb{R}^{2}:z_{1}>0\;\text{and}\;z_{2}<0\}\;.$
@@
$P_{2\sim 3}$ =$\;=\;\{z\in\mathbb{R}^{2}:z_{1}=0\;\text{and}\;z_{2}>0\}\qquad P_{3\sim 4}\;=\;\{z\in\mathbb{R}^{2}:z_{1}>0\;\text{and}\;z_{2}=0\}$ (8)
$P_{1\sim 2}$ =$\;=\;\{z\in\mathbb{R}^{2}:z_{1}<0\;\text{and}\;z_{2}=0\}\qquad P_{4\sim 1}\;=\;\{z\in\mathbb{R}^{2}:z_{1}=0\;\text{and}\;z_{2}<0\}\;.$
Finally let $\widehat{\mathcal{P}}=\cup_{i\in[4]}\widehat{P}_{i}$, where $\widehat{P}_{i}=P_{i}\cup P_{i\sim(i+1)}$. Observe by definition that $\widehat{\mathcal{P}}\cup\{(0,0)\}=\mathbb{R}^{2}$.
@@
[…]
@@
Due to the relationship between the coordinates of $y^{t}_{1}$ and $y^{t}_{2}$ from Proposition 4.2, the following relationship between $\Psi$ and $\psi$ (and by Proposition 3.4, between $\Psi$ and $\textnormal{Reg}(T)$), is immediate:
@@ 5 Discussion and Conclusion
In this work, we established for the first time that the unregularized Optimistic Fictitious Play algorithm can obtain constant $O(1)$ regret in two-player zero-sum games. Our proof technique leverages a geometric viewpoint of Fictitious Play algorithms, and we believe the techniques established for the 2$\times$2 regime can be extended to higher dimensions.
@@
Dimension: 15$\times$15 25$\times$25 50$\times$50
@@
Formally proving a constant regret bound for Optimistic Fictitious Play in all zero-sum games remains an important open question.
@@ Acknowledgements.
JL and RS were supported by the MOE Tier 2 Grant (MOE-T2EP20223-0018), the National Research Foundation, Singapore, under its QEP2.0 programme (NRF2021-QEP2-02-P05), the National Research Foundation Singapore and DSO National Laboratories under the AI Singapore Programme (Award Number: AISG2-RP-2020-016). SS was supported by the Villum Young Investigator Award (Grant no. 72091). The authors thank Anas Barakat and Andre Wibisono for helpful discussions.
@@ A.3 Details on Fictitious Play Variants as Skew-Gradient Descent
[…]
@@ Appendix B Assumptions on Payoff Matrix
Recall from Section 4 that in the proof of Theorem 3.5 for the 2$\times$2 setting, we make the following assumption on the entries of the payoff matrix:
@@
$\widetilde{A}=A+\Big{(}\frac{\det A}{a+b-(c+d)}\Big{)}\cdot\mathbf{1}\;.$
@@
Thus, by switching the maximization or minimization role between the players (via scaling the matrix by -1), we may assume that $a>0$. Finally, to show $a>\max\{b,c\}$ holds without loss of generality, observe from (17) that if $a+d-(b+c)>0$, then the interior Nash condition in (17) implies $a>c$ and $a>b$. If instead $a+d-(b+c)<0$, then $0<a<\min\{b,c\}$, and we can then rewrite the bilinear objective of the zero-sum game using a new payoff matrix with relabeled strategies (i.e., permuting the columns of $A$), as
@@
Under $A^{\prime}$, we have $b+c-(a+d)>0$, which from (17) and the reasoning above implies $b>\max\{a,d\}>0$. As a consequence, by possibly permuting the columns of $A$ and relabeling the strategies of Player 1, we can assume in either case that $a>\max\{b,c\}$. Together, we conclude that the assumption $a>\max\{0,b,c\}$ holds without loss of generality.
Similarly, it follows that we may also assume $d>\max\{0,b,c\}$ without loss of generality. Specifically, using the relabeling argument above, we may assume $a+d-(b+c)>0$. Then under the unique interior Nash and $\det A=0$ assumptions, it follows from (17) (using similar arguments as for $a\neq 0$) that $d\neq 0$ and $d>\max\{b,c\}$. Moreover, as $a>0$, if also $d<0$, then this implies $c,b<0$, meaning $\det A=ab-cd<0$, contradicting the assumption that $\det A=0$. Thus also $d>0$, and we conclude that the assumption $d>\max\{0,b,c\}$ holds without loss of generality. ∎
@@
[…]
Thus for all $t$, we have $\langle v_{1},y^{t}_{1}\rangle=(d-c)\cdot y^{t}_{11}+(a-b)\cdot y^{t}_{12}=0$ Rearranging, and recalling that $\rho_{1}:=(d-c)/(a-b)>0$ (where the inequality follows by Assumption 1), we find $y^{t}_{12}=-\rho_{1}\cdot y^{t}_{11}$.
@@
[…]
@@ Appendix C Proofs for Optimistic Fictitious Play Regret Upper Bound
In this section, we develop the omitted proofs from Section 4 that are needed to establish the main technical result of Theorem 3.5 (showing Optimistic FP has bounded energy in 2$\times$2 games).
@@ C.1 Properties of the Energy Threshold $B$
By definition of $B$, the level set =$\mathcal{L}=\{z\in R^{2}:\psi(z)=B\}$ must intersect the boundary of the ball =$\mathcal{B}=\{z\in\mathbb{R}^{2}:\|z\|_{1}\leq B\}$ on at least one of the boundaries $P_{i\sim(i+1)}$ in the set $\widetilde{\mathcal{P}}$ from Definition 4.3. Let $\mathcal{I}=\widetilde{\mathcal{P}}\cap\mathcal{B}\cap\mathcal{L}$ be the intersection of these three sets. Using the definition of $\psi$ from (10) it follows that for $z\in\mathcal{I}$
@@
Let $B$ be the constant from Definition 4.3. Suppose $\psi(z)>B$ and suppose $z\in\widehat{P}_{i}\cup P_{(i-1)~{}i}$ for some $i\in[4]$. Assume either $\widetilde{z}=z+S_{j}$ or $\widetilde{z}=z+S_{j}+S_{k}$ for $j,k\in[4]$ and $S$ as in (OFP Dual). Then
@@
The proof of Lemma C.2 proceeds using three separate propositions, which we state and prove as follows. First, we establish regions in which $\widetilde{z}^{t+1}$ and $z^{t+1}$ cannot lie when the energy $\psi$ is sufficiently large.
@@ D.1 Details on Alternating Play and Alternating Regret
  • (Initialization)  Assume without loss of generality $T$ is even. At time $t=1$, Player 1 chooses an initial $x^{1}_{1}\in\Delta_{m}$, and Player 2 observes $-A^{\top}x^{1}_{1}$.
@@
    Player 1 observes $\text{Player 1 observes }Ax^{t}_{2}\qquad\qquad\qquad\;\text{Player 2 observes }-A^{\top}x^{t-1}_{1}.$
@@
    Player 1 observes $\text{Player 1 observes }Ax^{t-1}_{2}\qquad\qquad\quad\;\text{Player 2 observes }-A^{\top}x^{t}_{1}.$
@@
[…]
@@ D.2 Details on Alternating Fictitious Play
$x^{t+1}_{1}$ $\;:=\;\mathop{\textnormal{argmax}}_{x\in\{e_{i}\}_{m}}\big{\langle}x,\sum_{k=1}^{t/2}A(x^{2k}_{2}+x^{2k-2}_{2})\big{\rangle}\quad\text{and}\quad x^{t+1}_{2}=x^{t}_{2}$
$x^{t+2}_{2}$ $\;:=\;\mathop{\textnormal{argmax}}_{x\in\{e_{i}\}_{n}}\big{\langle}x,\sum_{k=1}^{t/2}-A^{\top}(x^{2k+1}_{1}+x^{2k-1}_{2})\big{\rangle}\quad\text{and}\quad x^{t+2}_{1}=x^{t+1}_{1}\;.$
@@
($t$ even) $\begin{cases}x^{t}_{1}=x^{t-1}_{1}\\ x^{t}_{2}=\mathop{\textnormal{argmax}}_{x\in\{e_{i}\}_{n}}\;\langle x,y^{t}_{2}\rangle\end{cases}\quad\text{and}\quad\begin{cases}y^{t+1}_{1}=y^{t}_{1}+Ax^{t}_{2}\\ y^{t+1}_{2}=y^{t}_{2}-A^{\top}x^{t-1}_{1}\;.\end{cases}$ (AFP)
($t$ odd) $\begin{cases}x^{t}_{1}=\mathop{\textnormal{argmax}}_{x\in\{e_{i}\}_{m}}\;\langle x,y^{t}_{1}\rangle\\ x^{t}_{2}=x^{t-1}_{2}\end{cases}\quad\text{and}\quad\begin{cases}y^{t+1}_{1}=y^{t}_{1}+Ax^{t-1}_{2}\\ y^{t+1}_{2}=y^{t}_{2}-A^{\top}x^{t}_{1}\;.\end{cases}$
@@
$(t=3)\qquad\begin{cases}z^{3}=(-1,-2(2p-1))\in P_{1}\\ \widetilde{z}^{4}=(z^{3}_{1},z^{2}_{2})=(-1,-(2p-1))\in P_{1}\end{cases}\qquad w^{3}=\textnormal{Q}(\widetilde{z}^{4})=e_{1}$
$(t=4)\qquad\begin{cases}z^{4}=z^{3}+(-1,1)=(-2,-4p+3)\in P_{1}\\ \widetilde{z}^{5}=(z^{3}_{1},z^{4}_{2})=(-1,-4p+3)\in P_{1}\end{cases}\qquad w^{4}=\textnormal{Q}(\widetilde{z}^{5})=e_{1}$
$(t=5)\qquad\begin{cases}z^{5}=z^{4}+(-1,1)=(-3,-4p+4)\in P_{2}\\ \widetilde{z}^{6}=(z^{5}_{1},z^{4}_{2})=(-3,-4p+3)\in P_{1}\end{cases}\qquad w^{5}=\textnormal{Q}(\widetilde{z}^{6})=e_{1}$
$(t=6)\qquad\begin{cases}z^{6}=z^{5}+(-1,1)=(-4,-4p+5)\in P_{2}\\ \widetilde{z}^{7}=(z^{5}_{1},z^{6}_{2})=(-3,-4p+5)\in P_{2}\end{cases}\qquad w^{6}=\textnormal{Q}(\widetilde{z}^{7})=e_{2}.$
@@ E.1 Details on Experimental Setup
[…]
@@ E.2 Empirical Regret Comparisons of Fictitious Play and Optimistic Fictitious Play
Dimension: 15$\times$15 25$\times$25 50$\times$50

### v2

@@ 4.1 Subspace Dynamics of Optimistic Fictitious Play
$P_{2}$ $\;=\;\{z\in\mathbb{R}^{2}:z_{1}<0\;\text{and}\;z_{2}>0\}\hskip 18.49988ptP_{3}\;=\;\{z\in\mathbb{R}^{2}:z_{1}>0\;\text{and}\;z_{2}>0\}$ (7)
$P_{1}$ $\;=\;\{z\in\mathbb{R}^{2}:z_{1}<0\;\text{and}\;z_{2}<0\}\hskip 18.49988ptP_{4}\;=\;\{z\in\mathbb{R}^{2}:z_{1}>0\;\text{and}\;z_{2}<0\}\;.$
@@
$P_{2\sim 3}$ $\;=\;\{z\in\mathbb{R}^{2}:z_{1}=0\;\text{and}\;z_{2}>0\}\hskip 18.49988ptP_{3\sim 4}\;=\;\{z\in\mathbb{R}^{2}:z_{1}>0\;\text{and}\;z_{2}=0\}$ (8)
$P_{1\sim 2}$ $\;=\;\{z\in\mathbb{R}^{2}:z_{1}<0\;\text{and}\;z_{2}=0\}\hskip 18.49988ptP_{4\sim 1}\;=\;\{z\in\mathbb{R}^{2}:z_{1}=0\;\text{and}\;z_{2}<0\}\;.$
[…]
@@
Note that for notational convenience, when using an index $i\in[4]$ in the context of the sets of Definition 4.3, we assume addition and subtraction to $i$ are performed $(\textnormal{mod}\;4)$ in the natural way that maps to the set $\{1,2,3,4\}$. For example, $P_{i\sim(i+1)}$ is the set $P_{4\sim 1}$ when $i=4$, and $P_{i-2}=P_{i+2}$ is the set $P_{3}$ when $i=1$, etc. Figure 2 depicts the sets from Definition 4.3.
@@
Due to the relationship between the coordinates of $y^{t}_{1}$ and $y^{t}_{2}$ from Proposition 4.2, the following relationship between $\Psi$ and $\psi$ (and by Proposition 3.4, between $\Psi$ and $\textnormal{Reg}(T)$) is immediate:
@@ 5 Discussion and Conclusion
In this work, we established for the first time that the unregularized Optimistic Fictitious Play algorithm can obtain constant $O(1)$ regret in two-player zero-sum games. Our proof technique leverages a geometric viewpoint of Fictitious Play algorithms, and we believe the techniques established for the $\times2$ regime can be extended to higher dimensions.
@@
Dimension: $\times15$ $\times25$ $\times50$
@@
Limitations. Formally proving whether Optimistic Fictitious Play obtains constant regret in all zero-sum games remains an important open question.
Broader impact. We acknowledge that there are many potential societal consequences of our theoretical results, however none of which we feel must be specifically highlighted.
@@ Acknowledgements.
JL and RS were supported by the MOE Tier 2 Grant (MOE-T2EP20223-0018), the National Research Foundation, Singapore, under its QEP2.0 programme (NRF2021-QEP2-02-P05), the National Research Foundation Singapore and DSO National Laboratories under the AI Singapore Programme (Award Number: AISG2-RP-2020-016), and SS was supported by the Villum Young Investigator Award (Grant no. 72091). The authors thank Anas Barakat and Andre Wibisono for helpful discussions.
@@ A.3 Details on Fictitious Play Variants as Skew-Gradient Descent
$\;=\;\big\langle\partial\Psi(y^{t}),J\partial\Psi(\widetilde{y}^{t+1})\big\rangle\;=\;\big\langle\partial\Psi(y^{t}),J\partial\Psi(y^{t})\big\rangle\;=\;0\;.$
@@ Appendix B Assumptions on Payoff Matrix
Recall from Section 4 that in the proof of Theorem 3.5 for the $\times2$ setting, we make the following assumption on the entries of the payoff matrix:
@@
$\widetilde{A}=A-\Big(\frac{\det A}{a+d-(b+c)}\Big)\cdot\mathbf{1}\;.$
@@
Thus, by switching the maximization or minimization role between the players (via scaling the matrix by -1), we may assume that $a>0$. Finally, to show $a>\max\{b,c\}$ holds without loss of generality, observe from (17) that if a+d(b+c)>0a+d-(b+c)>0, then the interior Nash condition in (17) implies $a>c$ and $a>b$. If instead a+d(b+c)<0a+d-(b+c)<0, then $0<a<\min\{b,c\}$, and we can then rewrite the bilinear objective of the zero-sum game using a new payoff matrix with relabeled strategies (i.e., permuting the columns of $A$), as
@@
Under $A^{\prime}$, we have b+c(a+d)>0b+c-(a+d)>0, which from (17) and the reasoning above implies $b>\max\{a,d\}>0$. As a consequence, by possibly permuting the columns of $A$ and relabeling the strategies of Player 1, we can assume in either case that $a>\max\{b,c\}$. Together, we conclude that the assumption $a>\max\{0,b,c\}$ holds without loss of generality.
Similarly, it follows that we may also assume $d>\max\{0,b,c\}$ without loss of generality. Specifically, using the relabeling argument above, we may assume a+d(b+c)>0a+d-(b+c)>0. Then under the unique interior Nash and $\det A=0$ assumptions, it follows from (17) (using similar arguments as for $a\neq 0$) that $d\neq 0$ and $d>\max\{b,c\}$. Moreover, as $a>0$, if also $d<0$, then this implies $c,b<0$, meaning $\det A=ab-cd<0$, contradicting the assumption that $\det A=0$. Thus also $d>0$, and we conclude that the assumption $d>\max\{0,b,c\}$ holds without loss of generality. ∎
@@
$\langle v_{1},y^{t}_{1}\rangle\;=\;\sum_{k=1}^{t-1}\langle v_{1},Ax^{k}_{2}\rangle\;=\;0\;.$
Thus for all $t$, we have $\langle v_{1},y^{t}_{1}\rangle=(d-c)\cdot y^{t}_{11}+(a-b)\cdot y^{t}_{12}=0$ . Rearranging, and recalling that $\rho_{1}:=(d-c)/(a-b)>0$ (where the inequality follows by Assumption 1), we find $y^{t}_{12}=-\rho_{1}\cdot y^{t}_{11}$.
@@
$\langle v_{2},y^{t}_{2}\rangle\;=\;\sum_{k=1}^{t-1}\langle v_{2},-A^{\top}x^{k}_{1}\rangle\;=\;\sum_{k=1}^{t-1}\langle x^{k}_{1},-Av_{2}\rangle\;=\;0\;.$
@@ Appendix C Proofs for Optimistic Fictitious Play Regret Upper Bound
In this section, we develop the omitted proofs from Section 4 that are needed to establish the main technical result of Theorem 3.5 (showing Optimistic FP has bounded energy in $\times2$ games).
@@ C.1 Properties of the Energy Threshold $B$
By definition of $B$, the level set $\mathcal{L}=\{z\in R^{2}:\psi(z)=B\}$ must intersect the boundary of the ball $\mathcal{B}=\{z\in\mathbb{R}^{2}:\|z\|_{1}\leq B\}$ on at least one of the boundaries $P_{i\sim(i+1)}$ in the set $\widetilde{\mathcal{P}}$ from Definition 4.3. Let $\mathcal{I}=\widetilde{\mathcal{P}}\cap\mathcal{B}\cap\mathcal{L}$ be the intersection of these three sets. Using the definition of $\psi$ from (10) it follows that for $z\in\mathcal{I}$
@@
Let $B$ be the constant from Definition 4.3. Suppose $\psi(z)>B$ and suppose $z\in\widehat{P}_{i}\cup P_{(i-1)~i}$ for some $i\in[4]$. Assume either $\widetilde{z}=z+S_{j}$ or $\widetilde{z}=z+S_{j}+S_{k}$ for $j,k\in[4]$ and SS as in (OFP Dual). Then
@@
The proof of Lemma C.2 proceeds using three separate propositions, which we state and prove as follows:
First, we establish regions in which $\widetilde{z}^{t+1}$ and $z^{t+1}$ cannot lie when the energy $\psi$ is sufficiently large.
@@ D.1 Details on Alternating Play and Alternating Regret
  • (Initialization)  Assume without loss of generality $T$ is even. At time $t=1$, Player 1 chooses an initial $x^{1}_{1}\in\Delta_{m}$, and Player 2 observes $A^{\top}x^{1}_{1}$.
@@
    $\text{Player 1 observes }Ax^{t}_{2}\qquad\qquad\qquad\;\text{Player 2 observes }-A^{\top}x^{t-1}_{1}.$
@@
    $\text{Player 1 observes }Ax^{t-1}_{2}\qquad\qquad\quad\;\text{Player 2 observes }-A^{\top}x^{t}_{1}.$
@@
$\text{Player 1 cumulative payoff:}\quad\sum_{k=1}^{T/2}\big\langle x^{2k-1}_{1},A(x^{2k}_{2}+x^{2k-2}_{2})\big\rangle\;.$ (35)
$\text{Player 2 cumulative payoff:}\quad\sum_{k=1}^{T/2}\big\langle x^{2k}_{2},-A^{\top}(x^{2k+1}_{1}+x^{2k-1}_{1})\big\rangle\;.$
@@ D.2 Details on Alternating Fictitious Play
$x^{t+1}_{1}$ $\;=\;\mathop{\textnormal{argmax}}_{x\in\{e_{i}\}_{m}}\big\langle x,\sum_{k=1}^{t/2}A(x^{2k}_{2}+x^{2k-2}_{2})\big\rangle\quad\text{and}\quad x^{t+1}_{2}=x^{t}_{2}$
$x^{t+2}_{2}$ $\;=\;\mathop{\textnormal{argmax}}_{x\in\{e_{i}\}_{n}}\big\langle x,\sum_{k=1}^{t/2}-A^{\top}(x^{2k+1}_{1}+x^{2k-1}_{2})\big\rangle\quad\text{and}\quad x^{t+2}_{1}=x^{t+1}_{1}\;.$
@@
[…]
@@
$(t=3)\qquad\begin{cases}z^{3}=(-1,-2(2p-1))\in P_{1}\\ \widetilde{z}^{4}=(z^{3}_{1},z^{2}_{2})=(-1,-(2p-1))\in P_{1}\end{cases}$ $w^{3}=\textnormal{Q}(\widetilde{z}^{4})=e_{1}$
$(t=4)\qquad\begin{cases}z^{4}=z^{3}+(-1,1)=(-2,-4p+3)\in P_{1}\\ \widetilde{z}^{5}=(z^{3}_{1},z^{4}_{2})=(-1,-4p+3)\in P_{1}\end{cases}$ $w^{4}=\textnormal{Q}(\widetilde{z}^{5})=e_{1}$
$(t=5)\qquad\begin{cases}z^{5}=z^{4}+(-1,1)=(-3,-4p+4)\in P_{2}\\ \widetilde{z}^{6}=(z^{5}_{1},z^{4}_{2})=(-3,-4p+3)\in P_{1}\end{cases}$ $w^{5}=\textnormal{Q}(\widetilde{z}^{6})=e_{1}$
$(t=6)\qquad\begin{cases}z^{6}=z^{5}+(-1,1)=(-4,-4p+5)\in P_{2}\\ \widetilde{z}^{7}=(z^{5}_{1},z^{6}_{2})=(-3,-4p+5)\in P_{2}\end{cases}$ $w^{6}=\textnormal{Q}(\widetilde{z}^{7})=e_{2}.$
@@ E.1 Details on Experimental Setup
First, we note that all code used to run experiments can be found in the supplementary material. In this paper, all experiments were run locally on a single personal computer.
@@
To evaluate the robustness of regret guarantees to the tiebreaking method, we run the FP variants using both (a) fixed lexicographical tiebreaking (e.g., as in Abernethy et al., 2021a ) and (b) uniformly random tiebreaking (e.g., over the entries of the $\mathop{\textnormal{argmax}}$ set).
@@ E.2 Empirical Regret Comparisons of Fictitious Play and Optimistic Fictitious Play
Dimension: $\times15$ $\times25$ $\times50$